OpenStudy (anonymous):

30% of a number is 120. What is 75% of the number? A. 90 B. 300 C. 27 D. 48

OpenStudy (anonymous):

answer is b

OpenStudy (anonymous):

Use this formula my math teacher uses it helps me: \[\frac{ is }{ of } \times \frac{ % }{ 100 }\] Insert the numbers they give you into that formula and solve for x.

OpenStudy (anonymous):

It is supposed to be % over 100

OpenStudy (anonymous):

Lets put this into an equation, it says 30% of a number= 120 and we do not know a number that that is multiplied my 30% to get 120. so let X= a number then we would put the 30% into a decimal by dividing it by 100 which is 0.3 so the equation would be 0.3x=120 when you solve for x the answer is 400 and then multiply 0.75 by 400 to get the answer which is 300 so answer is B

OpenStudy (anonymous):

i think a

OpenStudy (anonymous):

its not A, Its B

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

tell me how u got A